# Robust Adaptive Discrete-Time Control Barrier Certificate

### v1

@@ 7 Conclusions
In this paper, we have developed a safe control method for discrete-time systems using robust adaptive discrete-time control barrier functions (CBFs), considering both parametric model uncertainty and disturbances. Unlike traditional approaches in continuous-time adaptive safe control design that tightly couple adaptation and control, our framework supports a modular architecture where the parameter learning algorithm operates independently of the safety filter. This separation simplifies safe control design and supports greater flexibility in implementation. In addition, the proposed method provides rigorous guarantees of safety without requiring exact model knowledge, which is particularly valuable for safety-critical applications where online adaptation is necessary. Future research directions include suboptimality analysis of safe control using CBFs and extension to distributed safe control of multi-agent systems.
@@ Appendix A Proof of Theorem 2
It suffices to prove robust safety through positive invariance. Considering $B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$, which characterizes the incremental value of the barrier function, it holds that
@@
$=$ $B(x_{t+1})-B(x_{t})-\frac{1}{2}(e_{\theta,t+1}\hskip-2.27626pt-\hskip-2.27626pte_{\theta,t})^{\top}\Gamma^{-1}(e_{\theta,t+1}\hskip-2.27626pt+\hskip-2.27626pte_{\theta,t})$
@@
$\geq$ $B(x_{t+1})-B(x_{t})-\frac{\|\delta_{\theta,t}\|^{2}+2\|\delta_{\theta,t}\|\varepsilon_{\theta,t}(p)}{2\lambda_{\min}(\Gamma)},\vskip-14.22636pt$ (28)
where $B(x_{t+1})-B(x_{t})$ can be lower bounded as
$B(x_{t+1})-B(x_{t})$
$=$ $B(f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t})-B(x_{t})$
[…]
By substituting the above equation into (A) and using the fact that $\pi_{t}(x_{t})$ satisfies (16), we have
$B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})\\ \geq-\alpha\left(B(x_{t})-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\geq-\alpha(B_{\mathrm{ra},t}(x_{t})).\vskip-5.69046pt$ (29)
Based on (29), if $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$, i.e., $B_{\mathrm{ra},t}(x_{t})\geq 0$, then $B_{\mathrm{ra},t+1}(x_{t+1})\geq B_{\mathrm{ra},t}(x_{t})-\alpha(B_{\mathrm{ra},t}(x_{t}))\geq 0$ since $\alpha\in\mathcal{C}_{\mathrm{SL}}$, which implies $x_{t+1}\in\mathcal{S}_{\mathrm{ra},t+1}$. Thus, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ is sequentially positive invariant for the closed-loop system (18) given $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$. Moreover, noting that $B_{\mathrm{ra},t}(x)\geq 0\Longrightarrow B(x)\geq 0$, (29) indicates that $\mathcal{S}_{B}$ is positive invariant for (18) over $\mathbb{I}_{[0:T]}$.       ∎
@@ Appendix B Robustified CBFs with error bounds
The true error $e_{\theta,t}$ is used to construct $B_{\mathrm{ra},t}$ in (14) to achieve adaptive safety, which is similar to [38]. Although safety can be established, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ cannot be characterized since $e_{\theta,t}$ is unknown. Considering that the ultimate goal is to achieve safety w.r.t. $\mathcal{S}_{B}$, which encodes safety [25], [38], the proposed method using $B_{\mathrm{ra},t}$ defined in (14) is still valid since $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B}$.
@@
The bound in (30) is conservative since it is solely computed using $\Theta_{t}$. Besides, the $\max$ ($\min$) operation prevents obtaining an analytical expression of its time derivative, and proving set invariance in continuous-time settings has to resort to the true error instead of the bound itself that is used to construct the CBF (see the proof of [25, Theorem 2]). The consequence is that, in the continuous-time setting, it is impossible [^5] Noting that, in the proof of [25, Theorem 2], $h\geq h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}$ and $h\geq 0$ cannot lead to $h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}\geq 0$. to establish the invariance property of the safe sets defined by the robustified CBFs constructed using the error bound. Luckily, in the discret-time setting, such invaraince property can be established since no derivative is required. Define the error-bound robust adaptive CBF by
$\overline{B}_{\mathrm{ra},t}(x):=B(x)-\frac{1}{2}\eta_{\theta,t}^{\top}\Gamma^{-1}\eta_{\theta,t},\vskip-5.69046pt$ (31)
where $\Gamma$ is a symmetric positive definite matrix. Similar to (15), the safe set corresponding to $\overline{B}_{\mathrm{ra},t}$ is given by
@@
Let Assumption 1 hold. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, A continuous function $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ on $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B\big{(}f(x,u;\hat{\theta}_{t})\big{)}-B(x)-L_{B}\overline{w}-E_{\eta,t}(x)\\ \geq-\alpha\left(B(x)-\frac{\|\eta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\forall t\in\mathbb{I}_{[0:T-1]}\vskip-5.69046pt$ (33)
where $E_{\eta,t}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B}\|\bm{\phi}(x)\|+\frac{\|\delta_{\eta,t}\|}{\lambda_{\min}(\Gamma)}\right)\eta_{\theta,t}+\frac{\|\delta_{\eta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}$ with $L_{B}$ being the Lipschitz constant in (9) and $\delta_{\eta,t}=\eta_{\theta,t+1}-\eta_{\theta,t}$. In this case, the condition (33) is called the error-bound robust adaptive CBC against disturbances and parametric model mismatch.
The certificate (33) also defines a set of safe inputs for $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ as
@@
Following Theorem 2, a modified safety guarantee that uses the certificate (33) is given as follows:
Theorem 4.
Let Assumption 1 hold and assume $x_{0}\in\overline{\mathcal{S}}_{\mathrm{ra},0}$. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, If $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for the system (6) and $\mathtt{EST}$ on $\{\overline{B}_{\mathrm{ra},t}\}^{T}_{t=0}$, then the set $\overline{\mathcal{U}}_{\mathrm{ras}}(x;\hat{\theta}_{t})$ is non-empty for all $x\in\mathcal{X}$ and all $t\in\mathbb{I}_{[0,T-1]}$. In addition, if a sequence of time-varying state-feedback control policies $\overline{\pi}_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\overline{\pi}_{t}(x)\in\overline{\mathcal{U}}_{\mathrm{ras}}(x;\hat{\theta}_{t})$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
@@
is robustly safe w.r.t. $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\mathcal{S}_{B}$ over $\mathbb{I}_{[0:T]}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
@@
The advantage of using $\xi_{\theta,t}$ to design the safety certificate is that the safe set $\overline{B}_{\mathrm{ra},t}$ can be explicitly determined. Moreover, it holds that $\eta_{\theta,t+1}\leq\eta_{\theta,t}$ if $\Theta_{t+1}\subseteq\Theta_{t}$, leading to $\overline{\mathcal{S}}_{\mathrm{ra},t}\subseteq\overline{\mathcal{S}}_{\mathrm{ra},t+1}$ (i.e., the safe set is gradually expanding).
@@ Appendix C More on Inherent Robustness
We first prove Theorem 3 using $V_{t}$ defined in (25). The proof is inspired by the proof of [30, Theorem B.13]. {pf} If there exists a $t_{0}\in\mathbb{N}$ such that $x_{t_{0}}\in\mathcal{S}_{\mathrm{ra},t_{0}}$, then $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ holds for all $t\in\mathbb{I}_{[t_{0}:\infty]}$ by Theorem 2 since $\{B_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is sequentially positive invariant for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$. Thus, in this case, $\lim_{t\to\infty}\mathds{1}(x_{t}\in\mathcal{S}_{\mathrm{ra},t})=1$. On the other hand, assume $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$; then we have
@@
where the last inequality is due to $B_{\mathrm{ra},t}(x_{t})<0$ for $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$ and $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$. Consider the sequence $\{V_{t}(x_{t})\}^{\infty}_{t=0}$; given that $V_{t}(x_{t})\geq 0$ and (36), by the monotone convergence theorem [31], there exists a $V^{\ast}\geq 0$ such that $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}$. Equivalently, $\forall\epsilon>0$, $\exists T^{\ast}(\epsilon)$ such that $\forall t\geq T^{\ast}(\epsilon)$, $V_{t}(x_{t})-V^{\ast}<\epsilon$. Next, we prove that $V^{\ast}=0$ by contradiction. Assume $V^{\ast}>0$, and let $\epsilon=-\alpha(-V^{\ast})$. Then, for any $t\geq T^{\ast}(-\alpha(-V^{\ast}))$, it holds that (i) $V^{\ast}\leq V_{t}(x_{t})<V^{\ast}-\alpha(-V^{\ast})$ and (ii) $V^{\ast}\leq V_{t+1}(x_{t+1})<V^{\ast}-\alpha(-V^{\ast})$, implying $V_{t+1}(x_{t+1})-V_{t}(x_{t})>\alpha(-V^{\ast})$. On the other hand, by (36), we have $V_{t+1}(x_{t+1})-V_{t}(x_{t})\leq\alpha(-V^{\ast})$, which leads to a contradiction. Therefore, $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}=0$, i.e., $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$.             ∎
@@
Let Assumption 1 hold. If $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ for $T=+\infty$ (cf. Definition B.1, then $x_{t}$ converges to $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ defined by the robust adaptive CBFs $\{\overline{B}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically (cf. Definition 5.1) on $\mathcal{X}$ for the system (6) with $u_{t}\in\overline{\mathcal{U}}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t})$.

### v2

@@ 7 Conclusions & Future work
In this paper, we have developed an adaptive safe control method for discrete-time systems using control barrier functions, considering both parametric model uncertainty and disturbances. Unlike traditional continuous-time adaptive safe control designs that tightly couple adaptation and control, our framework supports a modular architecture such that the parameter estimator operates independently of the safety filter. Future research directions include developing safe control methods for output-feedback systems with imperfect state measurements, extending the framework to systems with nonlinearly parameterized uncertainty, incorporating active exploration mechanisms for safe learning under model uncertainty, and characterizing the existence of CBFs and designing CBFs for uncertain systems.
@@ Appendix A Proof of Theorem 2
First, due to the definition of $\Delta B^{\star}(x,u)$ and $B^{\star}(x)$, the increment bound (12), and $\bar{\varepsilon}_{\theta}(p)\geq\varepsilon_{\theta,t}(p)$, it can be obtained that (15) is a sufficient condition for (17). Therefore, if $B$ is a robust adaptive DT-CBF that satisfies (15), then $\mathcal{U}_{\mathrm{ras},t}(x;p)$ is non-empty. Next, it suffices to prove robust safety through positive invariance. Considering $B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$, which characterizes the increment of the barrier function, it holds that
@@
$=$ $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$
$-\frac{1}{2}(e_{\theta,t+1}-e_{\theta,t})^{\top}\Gamma^{-1}(e_{\theta,t+1}+e_{\theta,t})$
@@
$\geq$ $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})-\frac{\|\delta_{\theta,t}\|^{2}+2\|\delta_{\theta,t}\|\varepsilon_{\theta,t}(p)}{2\lambda_{\min}(\Gamma)},\vskip-14.22636pt$ (28)
where $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$ can be further relaxed using the Lipschitz perturbation bound as
$B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$
$\hskip-12.91663pt\geq B(f(x_{t},\pi_{t}(x_{t});\hat{\theta}_{t}),\hat{\theta}_{t})-B(x_{t},\hat{\theta}_{t})$
$-L_{B,x}(\|\bm{\phi}(x_{t})\|\varepsilon_{\theta,t}(p)+\overline{w})-L_{B,\theta}\|\delta_{\theta,t}\|.\vskip-5.69046pt$ (29)
By substituting (A) into (A) and using the fact that $\pi_{t}(x_{t})$ satisfies (15), we have
$\hskip-8.61108ptB_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})\\ \hskip-17.22217pt\geq-\alpha\left(B(x_{t},\hat{\theta}_{t})\hskip-2.27626pt-\hskip-2.27626pt\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\geq-\alpha(B_{\mathrm{ra},t}(x_{t})).\vskip-5.69046pt$ (30)
Based on (30), if $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$, i.e., $B_{\mathrm{ra},t}(x_{t})\geq 0$, then $B_{\mathrm{ra},t+1}(x_{t+1})\geq B_{\mathrm{ra},t}(x_{t})-\alpha(B_{\mathrm{ra},t}(x_{t}))\geq 0$ since $\alpha\in\mathcal{C}_{\mathrm{SL}}$, which implies $x_{t+1}\in\mathcal{S}_{\mathrm{ra},t+1}$. Thus, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ is sequentially positive invariant for the closed-loop system (19) given $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras},t}(x;p)$. Moreover, noting that $B_{\mathrm{ra},t}(x)\geq 0\Longrightarrow B(x,\hat{\theta}_{t})\geq 0$, then $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ is positive invariant for (19).               ∎
@@ Appendix B Robustified CBFs with error bounds
The true error $e_{\theta,t}$ is used to construct $B_{\mathrm{ra},t}$ in (13) to achieve adaptive safety, which is similar to [39]. Although safety can be established, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ cannot be characterized since $e_{\theta,t}$ is unknown. Considering that the ultimate goal is to achieve safety w.r.t. $\mathcal{S}_{B,\hat{\theta}_{t}}$, which encodes safety [25], [39], the proposed method using $B_{\mathrm{ra},t}$ defined in (13) is still valid since $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B,\hat{\theta}_{t}}$.
@@
The bound in (31) is conservative since it is solely computed using $\Theta_{t}$. Besides, the $\max$ ($\min$) operation prevents obtaining an analytical expression of its time derivative, and proving set invariance in continuous-time settings has to resort to the true error instead of the bound itself that is used to construct the CBF (see the proof of [25, Theorem 2]). The consequence is that, in the continuous-time setting, it is impossible [^9] Noting that, in the proof of [25, Theorem 2], $h\geq h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}$ and $h\geq 0$ cannot lead to $h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}\geq 0$. to establish the invariance property of the safe sets defined by the robustified CBFs constructed using the error bound. Luckily, in the discret-time setting, such invaraince property can be established since no derivative is required. Define the error-bound robustified CBF as
$\overline{B}_{\mathrm{ra},t}(x):=B(x,\hat{\theta}_{t})-\frac{1}{2}\eta_{\theta,t}^{\top}\Gamma^{-1}\eta_{\theta,t},\vskip-5.69046pt$ (32)
where $\Gamma$ is a symmetric positive definite matrix as in (13), and the safe set corresponding to $\overline{B}_{\mathrm{ra},t}$ is given by
@@
Let Assumption 1 hold. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, A continuous function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$\Delta B(x,u)-L_{B,x}\overline{w}-E^{\star}_{\eta}(x)\\ \geq-\alpha\left(B^{\star}(x)-\frac{\|\eta_{\theta,0}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-5.69046pt$ (34)
where $\Delta B(x,u)$ and $B^{\star}(x)$ are defined as in Definition 3.2, and $E^{\star}_{\eta}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B,x}\|\bm{\phi}(x)\|+L_{B,\theta}\right)\|\eta_{\theta,0}\|$ with $L_{B,x}$ and $L_{B,\theta}$ given in Assumption 1. In this case, the condition (34) is called the error-bound robust adaptive CBC against disturbances and parametric model mismatch.
Following Theorem 2, a modified safety guarantee that uses Definition B.1 is given as follows:
Theorem 4.
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ with its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, then for the error-bound robust adaptive online input condition
$B\big(f(x,u;\hat{\theta}_{t}),\hat{\theta}_{t}\big)-B(x,\hat{\theta}_{t})-L_{B,x}\overline{w}\\ -E_{\eta,t}(x)\geq-\alpha\left(B(x,\hat{\theta}_{t})-\frac{\|\eta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-5.69046pt$ (35)
the set $\overline{\mathcal{U}}_{\mathrm{ras},t}(x)$ defined as
@@
is non-empty, $\forall x\in\mathcal{X}$ and $t\in\mathbb{I}_{[0,T-1]}$. The term $E_{\eta,t}(x)$ in (17) equals $L_{B,x}\|\bm{\phi}(x)\|\|\eta_{\theta,t}\|+L_{B,\theta}\|\delta_{\eta,t}\|$ with $\delta_{\eta,t}=\eta_{\theta,t+1}-\eta_{\theta,t}$. In addition, if time-varying state-feedback control policies $\overline{\pi}_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\overline{\pi}_{t}(x)\in\overline{\mathcal{U}}_{\mathrm{ras},t}(x)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
@@
is robustly safe w.r.t. $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
@@
The advantage of using $\eta_{\theta,t}$ as in (32) is that $\overline{B}_{\mathrm{ra},t}$ can be explicitly determined. Moreover, using set-membership identification, it holds that $\eta_{\theta,t+1}\leq\eta_{\theta,t}$ (i.e., the error decays monotonically) since $\Theta_{t+1}\subseteq\Theta_{t}$, leading to $\overline{\mathcal{S}}_{\mathrm{ra},t}\subseteq\overline{\mathcal{S}}_{\mathrm{ra},t+1}$, i.e., the robustified safe set is monotonically expanding.
@@ Appendix C More on Inherent Robustness
We first prove Theorem 3 using $V_{t}$ defined in (25). The proof is inspired by the proof of [32, Theorem B.13].
{pf}
If there exists a $t_{0}\in\mathbb{N}$ such that $x_{t_{0}}\in\mathcal{S}_{\mathrm{ra},t_{0}}$, then $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ holds for all $t\in\mathbb{I}_{[t_{0}:\infty]}$ by Theorem 2 since $\{B_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is sequentially positive invariant for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$. Thus, in this case, $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$. On the other hand, assume $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$; then we have
@@
where the last inequality is due to $B_{\mathrm{ra},t}(x_{t})<0$ for $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$ and $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$. Consider the sequence $\{V_{t}(x_{t})\}^{\infty}_{t=0}$; given that $V_{t}(x_{t})\geq 0$ and (38), by the monotone convergence theorem [33], there exists a $V^{\ast}\geq 0$ such that $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}$. Equivalently, $\forall\epsilon>0$, $\exists T^{\ast}(\epsilon)$ such that $\forall t\geq T^{\ast}(\epsilon)$, $V_{t}(x_{t})-V^{\ast}<\epsilon$. Next, we prove that $V^{\ast}=0$ by contradiction. Assume $V^{\ast}>0$, and let $\epsilon=-\alpha(-V^{\ast})$. Then, for any $t\geq T^{\ast}(-\alpha(-V^{\ast}))$, it holds that (i) $V^{\ast}\leq V_{t}(x_{t})<V^{\ast}-\alpha(-V^{\ast})$ and (ii) $V^{\ast}\leq V_{t+1}(x_{t+1})<V^{\ast}-\alpha(-V^{\ast})$, implying $V_{t+1}(x_{t+1})-V_{t}(x_{t})>\alpha(-V^{\ast})$. On the other hand, by (38), we have $V_{t+1}(x_{t+1})-V_{t}(x_{t})\leq\alpha(-V^{\ast})$, which leads to a contradiction. Therefore, $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}=0$, implying $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$ by the continuity of $V_{t}$.   ∎
@@
Let Assumption 1 hold. If $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ (cf. Definition B.1), then $x_{t}$ converges to $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically (cf. Definition 5.1) on $\mathcal{X}$ for the system (6) with $u_{t}\in\overline{\mathcal{U}}_{\mathrm{ras},t}(x_{t})$.
@@
Biographies
Changrui Liu received the BSc degree (cum laude) from Harbin Institute of Technology (HIT) in Harbin, China, he graduated from Honors School, HIT and was awarded Best 10 Graduate, 2019. He then received the MSc degree from TU Delft, Aerospace Engineering. He is currently pursuing the PhD degree at the Delft Center for Systems and Control, TU Delft. His research interests are performance guarantees in machine learning and optimization.
Anil Alan received the B.Sc. degree from Mechanical Engineering Departments of Middle East Technical University, Ankara, Türkiye, in 2012, the M.Sc. degree from Bilkent University, Ankara, in 2017, and the Ph.D. degree from the University of Michigan, Ann Arbor, MI USA, in 2024. He is currently a Postdoc with the Delft University of Technology, Delft, The Netherlands. His research focuses on nonlinear safety-critical control with application to connected automated vehicles.
Shengling Shi is an assistant professor at Delft Center for Systems and Control, TU Delft, and he was a postdoctoral associate at the Massachusetts Institute of Technology, USA. He was also a postdoc at the Delft Center for Systems and Control, TU Delft, the Netherlands. He received the Ph.D. degree from the Eindhoven University of Technology in 2021. His research interests include system identification, model predictive control, and their applications.
Bart De Schutter (IEEE member since 2008, senior member since 2010, and fellow since 2019) is a full professor and head of department at the Delft Center for Systems and Control of Delft University of Technology in Delft, The Netherlands. He is also an IFAC fellow.